OpenStudy (anonymous):

A stadium has 52,000 seats. Seats sell for $42 in section A, $36 in section B, and $30 in section C. The number of seats in section A equals the total number of seats in sections B and C. Suppose the stadium takes in $1,960,200 from each sold out event. How many seats does each section hold?

OpenStudy (ddcamp):

Total seats = A + B + C Total sales = 42(A) + 36(B) + 30(C) A = B + C Solve the system of equations.

OpenStudy (anonymous):

I dont follow w how to set it up

OpenStudy (ddcamp):

A is the number of seats in section A B is the number of seats in section B C is the number of seats in section C There are 52000 seats total, so 52000 = A+B+C There are the same number of seats in section A as in B and C combined, so A = B + C Using the ticket price for seats in each area, and the sales when all the seats are full: 1960200 = 42(A) + 36(B) + 30(C)

OpenStudy (ddcamp):

Easiest start is: 52000 = A + (B+C); A = B+C → 52000 = 2A → 26000 = A With that, finding B and C should be fairly straightforward.

OpenStudy (anonymous):

Thanks, trying this now.

OpenStudy (anonymous):

total sales = 42(B+C) + 36B + 30C so $1,960,200 = 42B + 42C + 36B + 30C or 78B + 72C

OpenStudy (anonymous):

Confused.

OpenStudy (anonymous):

B=A-26000 is as far as I've gotten lol

OpenStudy (anonymous):

B=C-2600

OpenStudy (anonymous):

B=C-26000

OpenStudy (anonymous):

C+B+26000=52000